OpenStudy (anonymous):

Please help with Physics Practice question. The moon is about 3.8 x 10^5km from Earth. Show that its average orbital speed about Earth is 1026 m/s

OpenStudy (anonymous):

hmm.. i agree hold on let me see

OpenStudy (anonymous):

g (gravity) = G*M/R^2 a (centripetal) = V^2/R G*M/R^2 = V^2/R v = sqrt[G*M/R] Using M = 5.9742E+24 kg and G = 6.673E-11 And don't forget that the radius should be in meters and not km so R = 3.8E8 We get v = 1024.3 m/s You can also do a little more and get v in terms of the radius of the earth At the surface of the earth: g = G*M/r^2 which gives M = g*r^2/G v = sqrt[g*r^2/R] Using r = 6378.1 km gives v = 1024.8 m/s

OpenStudy (anonymous):

that really helped put it into perspective thank you